OpenStudy (anonymous):

Find the quotient. find the quotient (4ab - 8a 2 + 12ac) ÷ 4a

OpenStudy (dls):

\[\Huge \frac{4ab}{4a}-\frac{8a^2}{4a}-\frac{12ac}{4a}\]

OpenStudy (dls):

cut,cut and cat

OpenStudy (anonymous):

would it be b - 2a + 3c

OpenStudy (dls):

yup